### v3

# Trust-Oriented Adaptive Guardrails for Large Language Models

Abstract
Guardrail, an emerging mechanism designed to ensure that large language models (LLMs) align with human values by moderating harmful or toxic responses, requires a sociotechnical approach in their design. This paper addresses a critical issue: existing guardrails lack a well-founded methodology to accommodate the diverse needs of different user groups, particularly concerning access rights. Supported by trust modeling (primarily on ‘social’ aspect) and enhanced with online in-context learning via retrieval-augmented generation (on ‘technical’ aspect), we introduce an adaptive guardrail mechanism, to dynamically moderate access to sensitive content based on user trust metrics. User trust metrics, defined as a novel combination of direct interaction trust and authority-verified trust, enable the system to precisely tailor the strictness of content moderation by aligning with the user’s credibility and the specific context of their inquiries. Our empirical evaluation demonstrates the effectiveness of the adaptive guardrail in meeting diverse user needs, outperforming existing guardrails while securing sensitive information and precisely managing potentially hazardous content through a context-aware knowledge base. To the best of our knowledge, this work is the first to introduce trust-oriented concept into a guardrail system, offering a scalable solution that enriches the discourse on ethical deployment for next-generation LLM service.
[…]
6705
1 Introduction
Large language models (LLMs), such as GPT [1] and Llama [37], have significantly expanded the reach of Artificial Intelligence (AI) technologies across critical sectors such as healthcare, finance, and engineering [38]. However, with these advancements come significant challenges, particularly concerning the reliability, safety, robustness, and ethical use of such models [16]. Prominent issues such as hallucination, biases, and toxicity highlight the critical need for robust safeguards, a.k.a. guardrails, to ensure that LLMs operate within acceptable boundaries, and it has been argued that the design of such guardrails need to be sociotechnical in their design, considering both ‘social’ and ‘technical’ aspects [5, 7]. To this end, this paper revisits the trust modeling [41, 4] in the context of LLM safeguarding, and integrates it into the guardrail design in order to boost the guardrail’s ability to cater for different needs from a diverse set of users.
In addition to the above ‘social’ need, this work also addresses a ‘technical’ concern that existing guardrails are rule-based [31, 17] or incorporating extensive post-processing checks [42] in LLMs applications. While these approaches offer some level of control, they are often static, inflexible and unable to adapt to the dynamic nature of human activities. As a result, there is a growing need for more adaptive and context-sensitive solutions that can better manage the risks associated with deploying LLMs in real-world scenarios.
Consider a scenario depicted in Figure 1a, a Policeman who is attempting to access criminal design information can be hindered by static guardrails. Despite the Policeman’s legitimate professional need, the static guardrails’ stringent and inflexible parameters deny access to all sensitive information. The repeated access denials persist even when attempts are made to manually relax system constraints for trustable user, compromising both practicality and user experience of LLMs. This rigidity often prevents specialized users from accessing essential information, as demonstrated when all of the Policeman’s legitimate requests are denied despite verified credentials. This situation highlights the limitations of static guardrails which are unable to gradually granting access to the specialized sensitive knowledge through evaluating the policeman’s credentials, job role, consistent interaction history and other relevant factors.
Figure 1: Comparison of Static vs. Adaptive Guardrails in Managing Access to Domain-specific Sensitive Information
This paper explores trust modeling, a mechanism that is perceived as the basis for decision making in many contexts for maintaining long-term relationships based on cooperation and collaboration [4], to address these issues. In other safety-critical areas, trust is utilized to adapt systems’ capabilities, ensuring consistent, safe, and efficient decisions when interacting with users [28, 40]. Nevertheless, directly implementing trust modeling in LLMs’ guardrails faces several challenges: 1) user profiles are normally private for LLM applications; 2) the trust dimension is difficult to quantify across various application contexts; 3) Difficulty exists in accurately assessing trust during dynamic interactions while balancing risk with data accessibility. In addition to trust modeling, In-context Learning (ICL) via Retrieval-Augmented Generation (RAG) plays a crucial role in enhancing the adaptability of guardrails. It empowers models to swiftly adapt their outputs without updating parameters, guided by specific instructions or contextual examples presented by trust-oriented retrieved context [11]. Leveraging this capability, it is possible to implement guardrails that are not only adaptive but also finely responsive to the sensitive queries encountered by the LLM, guided by hierarchical knowledge sources and aligned with user trust levels. This synergy between trust modeling and ICL offers a novel framework for ensuring the safety of LLMs while maintaining practicality.
In this paper, we propose an adaptive guardrails mechanism illustrated in Figure 1b for LLMs, which dynamically adjusts content moderation and information richness based on individual user trust modeling. Unlike static guardrails, this adaptive system allows users such as a policeman to access sensitive information by presenting relevant trust attributes. These are then integrated into a composite trust score calculated from both $Direct\ Interaction\ Trust$ [^1] Derived from the user’s historical interactions with the LLM and $Authority\ Verified\ Trust$ [^2] Established via credential verification by Trusted Third Parties to determine the access level. Additionally, we use ICL to customize responses to highly sensitive queries according to the obtained trust score. The key contribution of this paper include the development of a trust model tailored to LLMs, the implementation of adaptive guardrails that adjust access level and information richness based on trust scores and ICL, along with empirical analysis of the effectiveness of our proposed approach.
2 Related Work
2.1 Existing Guardrails
Guardrails for LLMs are programmable, rule-based systems acting as intermediaries to ensure that interactions between users and models adhere to established ethical and operational standards [8]. Reflecting the latest advancement, several existing techniques represent the state of the art in this field. Llama Guard [17] operates as a binary classifier that effectively differentiates between safe and unsafe content, ensuring moderation aligns with predefined safety guidelines. Concurrently, Nvidia NeMo [31] enhances interaction accuracy by utilizing the K-nearest neighbors method to align user intents with the most appropriate vector-based canonical forms. After this matching process, LLMs generate safe and pertinent responses, guided by Colang scripts. Moreover, Guardrails AI [30] bolsters system reliability and trust by implementing structured safety protocols, involving detailed XML specifications that rigorously check that outputs adhere to strict content integrity and compliance standards. Supporting these technologies, Python packages such as LangChain [^3] https://www.langchain.com/, AI Fairness 360444https://github.com/Trusted-AI/AIF360, Adversarial Robustness Toolbox [^5] https://github.com/Trusted-AI/adversarial-robustness-toolbox, Fairlearn [^6] https://github.com/fairlearn/fairlearn, and Detoxify [^7] https://github.com/unitaryai/detoxify address biases, boost robustness, and ensure content safety and fairness, collectively fostering responsible and reliable LLM applications.
2.2 Trust Models
Trust is defined by the Merriam-Webster Dictionary [6] as “assured reliance on the character, ability, strength, or truth of someone or something," which succinctly encapsulates trust as a fundamental aspect of relationships, where a trustor places confidence in a trustee based on established criteria. Recognized as a pivotal element in decision-making, trust spans various disciplines including management, psychology, economics, and engineering, underscoring its multidisciplinary importance [21].
Recently, trust in human-technology relationships has gained widespread attention and is typically classified into three distinct types: credentials-based, cognitive trust, and experience-based [15]. Credentials-based trust is commonly employed in security systems, where users are required to provide valid credentials to verify their identity and gain access to services, satisfying established security policies [25]. Cognitive trust involves in the human aspects of trust, focusing on subjective judgments and trust-based decisions in interactions, especially between humans and robots, highlighting the psychological dimensions of trust [9]. Experience-based trust is frequently used in multiagent systems and e-service platforms, where a trustee’s reliability is assessed based on past interactions and reputation-based trust, helping to predict the trustees’ future behaviors [13, 33]. This approach helps predict future behaviors by employing statistical methods to calculate trust scores. In this paper, we introduce a composite trust model that identifies relevant dimensions from various trust paradigms to meet the dual needs of security and practicality in LLM-user interactions, comprehensively assessing trustee reliability in LLMs equipped with adaptive guardrails.
2.3 In-Context Learning
In-context learning utilizes the generalization capabilities of LLMs to efficiently perform tasks with just a few contextual examples, requiring minimal data and no parameter updates, setting it apart from standard in-weights learning like gradient-based fine-tuning [20]. This method has become a focal point due to its ability to streamline the learning process and reduce the computational overhead typically associated with training large models. Research has focused on the selection of demonstration examples and the better ways of problem formulations [44, 23]. A common enhancement to ICL is RAG, which retrieves external knowledge to refine LLM outputs with contextually rich information, improving accuracy and adaptability in real-world scenarios [35, 11]. In our paper, we extend this approach by leveraging RAG to implement online ICL, dynamically retrieving trust-oriented information from hierarchical knowledge bases. By incorporating trust-based criteria, our method ensures that the retrieved knowledge is not only contextually rich but also aligned with user-specific constraints, enabling LLMs to interact effectively with external agents or access expert-curated knowledge sources.
3 Problem Formulation
Existing guardrails, formally represented as $G$, often enforce uniform and rigid rules to LLMs’ response on $x$ against a set of predefined safety conditions $C$. This indiscriminate approach to safety may lead to overly restrictive systems that implement moderated responses $M(x)$, such as “I am sorry…", which do not adequately balance safety with usability and flexibility for diverse user groups.
$\small G(x)=\begin{cases}x&\text{if }x\models C\\ M(x)&\text{if }x\not\models C\end{cases}$ (1)
To overcome the limitations, we define several components which collaboratively adjust the strictness of content moderation based on user trust scores. Let $U$ be a set of users. For any user $u_{i}\in U$ with $i$ being the user index, a trust score $T_{i}$ is assigned. By considering adaptable trust scores, we refine content moderation depending on both predefined safety conditions $C$ and user-specific trust score $T_{i}$. Formally, the function of adaptive guardrails $AG$ is defined below:
$\small AG(x,T_{i})=\begin{cases}R(x,T_{i})&\text{if }x\models C\text{ or }(x\not\models C\text{ and }T_{i}\geq\beta)\\ M(x)&\text{if }x\not\models C\text{ and }T_{i}<\beta\end{cases}$ (2)
where $\beta$, initially assigned with a unified base value, serves as a trust threshold dynamically adjusted for different user groups and queries. Intuitively, for users verified by reliable authorities, $\beta$ is relaxed based on authority levels, while for less reliable authorities, $\beta$ is elevated according to query sensitivity to prevent misuse. $R(\cdot)$ is the content richness function, defined in Equation (14) to determine the level of sensitive information provided by knowledge base with hierarchal confidentiality. From Equation (2), it is clear that we keep the strict forbidden for users with low trust levels, but release the restriction for trustworthy user groups to ensure personalized access.
4 Methodology
In this section, we introduce an adaptive trust model that calculates trust score $T$, and through a contextual adaptation mechanism, utilizes $T$ to adjust content richness $R$ within moderation guardrails for LLMs. Departing from the traditional one-size-fits-all approach, our adaptive system involves a trust evaluation mechanism for potentially unsafe messages, as illustrated in Figure 2.
Inspired by Cho et al.’s work [4], which emphasizes the use of weighted aggregation to construct multi-dimensional trust in human-machine interactions, we adopt their classification of trust into communication, information, social, and cognitive dimensions to derive a composite trust score. Each dimension includes multiple trust attributes, from which we selected the most applicable ones for the adaptive guardrail context, with the flexibility to integrate others whenever needed. We design two modules following [29], i.e., Direct Interaction Trust ($DT$) and Authority-Verified Trust ($AT$). $DT$ is crucial in environments where advisors are dishonest or exhibit frequently changing behaviors, as it captures communication and information trust through user interaction history and content assessments. In contrast, $AT$ as indirect trust is more effective in contexts with reliable advisors, as it evaluates social and cognitive trust using third-party verification. By integrating real-time direct trust assessments with robust external validations, these modules are essential for ensuring the system remains adaptive and reliable [19]. Eventually, if the composite trust score calculated by these two modules surpasses a predefined threshold, LLMs can adaptively adjust their responses to enhance system integrity; otherwise, access is restricted to prevent potential misuse.
Figure 2: Trust Evaluation for Sensitive Information
The Composite Trust Rating ($T$), synthesized by $DT$ and $AT$, scaled between $[0,1]$ and utilizing principles of beta distribution for a probabilistic trust assessment. This rating reflects the level of guardrail restrictions, enabling precise adjustments to match user trust profiles. Higher $T$ values result in fewer restrictions, allowing the guardrails to align closely with demonstrated trust levels. The formula $T_{i}$ for user $u_{i}$ is presented as follows:
$\small T_{i}=\eta\cdot AT_{i}+(1-\eta)\cdot DT_{i}$ (3)
where $\eta$ is an adaptive weight that adjusts based on the authority rankings of Trust Third Parties (TTPs) and the threshold of direct interaction trust, $\delta$. This mechanism highlights that in environments with reliable TTPs, indirect trust can be more beneficial than direct trust [29]. On the other hand, when the assurance from TTPs is less reliable, the guardrails augment the $DT$ weight to ensure the security and integrity of LLMs’ inputs and responses, effectively managing access to sensitive information during dynamic interaction. The adaptive weight is formulated as below:
$\eta=\begin{cases}1&\mathcal{A}.ranking=top\\ \theta+\frac{1-\theta}{1+e^{(-I\cdot(DT_{i}-\delta))}}&\overline{DT_{i}}\geq\delta\land\mathcal{A}.ranking!=top\\ 0&\overline{DT_{i}}<\delta\lor\mathcal{A}.ranking=low\end{cases}$ (4)
In Equation (4), $\overline{DT}$ is the average direct interaction trust during the dialogue. $\theta$ serves as a scaling factor that adjusts the magnitude of trust modification and determines the baseline influence of $AT$, while $I$ acts as a regulator to affect the steepness of decay rate.
4.1 Direct Interaction Trust
Based on [36, 34, 12], Direct Interaction Trust should satisfy the following properties: 1) Trust gradually decreases over time, exhibiting a temporal correlation that strengthens as it approaches the present moment. 2) Additionally, trust should be strongly correlated with the consistency of interactions, and trust is more readily destroyed by negative experiences. To achieve them, we formulate a mathematical expression in Eq. (9), related to time and consistency during interactions, which is used to quantitatively compute the DT.
We introduce the time decay factor, as defined in Equation 5, which intuitively indicates how a target attribute is readily substituted by one that is more easily recalled. This substitution prioritizes recent interactions, emphasizing the greater significance of the most recent events.
$d(\tau,\tau_{t})=e^{-\gamma(\tau-\tau_{t})}$ (5)
where $\gamma$ is the decay constant, $\tau$ is the current time, and $\tau_{t}$ represents the time of interaction $t$. As $\tau-\tau_{t}$ increases, $d$ decreases, reducing the influence of older interactions.
Simultaneously, the trust model utilizes a dynamic update mechanism where each interaction updates the safe ($a$) and unsafe ($b$) message counts based on the time decay factor and sliding window mechanism, ensuring that only the most recent interactions significantly impact trust assessments:
$a^{(\tau)}_{i}$ $=\sum_{t=1}^{W}d(\tau,\tau_{t})\cdot a^{(\tau_{t})}_{i}+safe^{(\tau)}$ (6)
$b^{(\tau)}_{i}$ $=\sum_{t=1}^{W}d(\tau,\tau_{t})\cdot b^{(\tau_{t})}_{i}+unsafe^{(\tau)}$ (7)
where $a^{(\tau_{t})}_{i}$ and $b^{(\tau_{t})}_{i}$ represent the counts of safe and unsafe interactions at each past interaction $t$ within the sliding window $W$, which limits the summation to the most recent $W$ interactions. $safe^{(\tau)}$ and $unsafe^{(\tau)}$ are the corresponding counts at the current time $\tau$.
Moreover, in interactions involving LLMs, consistency always serves as a critical factor of interaction uniformity. It quantifies the alignment of a current interaction with historical interactions, providing insights into user behavior patterns. We define the interaction consistency score by applying a quadratic transformation to the cosine similarity metric, which compares embedding vectors from current and past interactions [43]. This transformation effectively scales the values to remain within the $[0,1]$ interval, while enhancing their directional alignment:
$\small IC^{(\tau)}_{i}=\frac{1}{W}\sum_{t=1}^{W}\left(\frac{1+\frac{\mathbf{v}\cdot\mathbf{v}_{t}}{\|\mathbf{v}\|\|\mathbf{v}_{t}\|}}{2}\right)^{2}$ (8)
where $\mathbf{v}$ is the feature vector representing the current interaction. $\mathbf{v}_{t}$ represents the feature vector of the $t$-th historical interaction within the sliding window. Finally, based on the above ingredients, we adopt Bayesian inference formula based on Beta probability density function to evaluate the direct interaction trust, inspired by [18, 3]. The interaction history and consistency score are combined to form direct interaction trust to offer a complete perspective for evaluating continuous and dynamic user engagement:
$DT_{i}^{(\tau)}=\frac{a^{(\tau)}_{i}+IC^{(\tau)}_{i}\cdot w+1}{a^{(\tau)}_{i}+b^{(\tau)}_{i}\cdot n+2}$ (9)
where $DT_{i}$ is the Direct Interaction Trust score for $u_{i}$ at time $\tau$, $w$ is the weighting factor for $IC$, and $n$ denotes the unsafe coefficient to amplify the impact of unsafe interactions.
4.2 Authority-Verified Trust
Authority-verified trust utilizes demonstrations from Trust Third Parties (TTPs). Inspired by prior research on recommendation systems [3], we introduce a trust algorithm and reinterpret it to address the unique needs of LLM guardrails, which incorporates similarity of views ($S$) and confidence level ($C$) as shown in Eq. (12). Specifically, $S$ is based on the principle that a trustor is more likely to accept suggestions from TTPs whose perspectives resonate more closely with its own evaluations [10]. In our context, we assess $S$ by comparing the historical average direct interaction trust and the normalized rating credits ($NR$) assigned by TTPs to corresponding trustees.
Additionally, confidence level $C$ is also a significant factor that should be considered in trust dynamics, as trust relies on the stability and reliability of expectations, which are determined by the level of confidence [24, 39]. It can be gauged by the volume of trust attributes that TTPs attribute to trustees, where a higher positive count indicates greater trust. Trust attributes encompass factors such as social reputation, job title, historical credibility, and family background, alongside other pertinent factors like professional achievements and peer evaluations. They present a comprehensive view of a user’s reliability and standing within their social networks. For the relevance and timeliness of these attributes, they are periodically refreshed through updated data requests from TTPs. The accumulated positive ($pos_{pu}$) and negative ($neg_{pu}$) trust attributes from a TTP $p_{k}$ about the $u_{i}$ are represented statistically and responsible for calculating TTPs’ confidence $C$, allowing quantify trust based on the diversity of trust attributes $j$.
$pos_{ki}^{(\tau)}$ $=\mathcal{A}\cdot\sum_{j=1}^{J}pos_{ki}^{j},\quad neg_{ki}^{(\tau)}$ $=\mathcal{A}\cdot\sum_{j=1}^{J}neg_{ki}^{j}$ (10)
To enhance the precision of area relevance ($AR$) assessments within our trust model, we employ Prompt Engineering to direct the inference process of multiple LLMs debating $\mathcal{D}$. The robustness of our final decision is strengthened by integrating semantic similarity measures, denoted as $ST$. Specifically, we employ the all-mpnet-base-v2 sentence transformer in our cases, which yields semantically rich embeddings that effectively capture nuanced linguistic relations [32]. This method refines the evaluation of area relevance between a trustee’s professional area, as verified by TTPs, and the sensitivity required by the prompts. The formula for computing $AR$ in $[0,1]$ is presented as follows:
$AR^{(\tau)}_{ki}=\frac{ST^{(\tau)}(Area,Input)+\mathcal{D}^{(\tau)}(Area,Input)}{2}$ (11)
Our model eventually integrates the authority indicator ($\mathcal{A}$), set by the guardrail system for TTPs, along with previously mentioned variables, to collectively ensure the authority-verified trust aligns with the practical and safety demands of the guardrail system. The definitive $AT$ formula for a trustee $u_{i}$ and $K$ TTPs, is designed to encapsulate these trust factors.
$AT_{i}^{(\tau)}=\sum_{k=1}^{K}\frac{\mathcal{A}_{k}^{(\tau)}S_{k}^{(\tau)}C_{ki}^{(\tau)}NR_{ki}^{(\tau)}AR_{ki}^{(\tau)}}{\sum_{k=1}^{K}\mathcal{A}_{k}^{(\tau)}S_{k}^{(\tau)}C_{ki}^{(\tau)}}$ (12)
In this paper, we consider a set of TTPs $p_{k}\in P$, where $k$ is the TTP index and these TTPs are divided by three ranking levels $\mathcal{A}.ranking\in\{top,medium,low\}$. Each level is assigned a weight indicator $\mathcal{A}\in[0,1]$, which quantifies their relative authority. Furthermore, TTPs are independent entities that provide verification services for systems, ensuring the authenticity and accuracy of data without direct involvement in the transactions or interactions between primary parties. For instance, validating a university email to confirm educational qualifications illustrates the necessity of TTPs in accessing specialized information for research purposes.
Remark 1: We also incorporate periodic re-validation throughout the whole trust evaluation process, which is a common practice in web services [2], to ensure that all sensitive information requests meet specific security requirements within guardrail systems. This mandatory procedure applies universally, irrespective of trust scores, preventing outdated or overly trusted credentials from granting inappropriate access and thus enhancing the system’s robustness. For long-term access to sensitive or controversial information for legal or specialized purposes, users must periodically re-validate their identity with TTPs to maintain credential validity.
4.3 Contextual Adaptation Mechanism
Upon calculating user trust scores, we classify adaptive guardrails into tiers that regulate control mechanisms. These tiers manage users’ access to sensitive information. Key hyperparameters of LLMs, such as temperature, response length, and token limits, are dynamically adjusted based on the user’s trust score, ensuring safe operation while accommodating personalization and trust. For example, users with higher trust scores may have higher token limits for generating responses when accessing sensitive information.
Figure 3: Contextual Adaptation via RAG and ICL
Furthermore, integrating RAG and ICL significantly enhances the relevance and specificity of responses in language models. In our system, RAG leverages an expertly constructed, hierarchically structured confidential knowledge base to retrieve trust-oriented contexts. This integration ensures that the response generation process is dynamically tailored according to both the contexts and the user’s trust score. The confidentiality of the access level is governed by $T_{i}$ and is defined as follows:
$AL_{i}=\sum_{l=1}^{L}\mathbf{1}(T_{i}\geq\xi_{l})$ (13)
where $\xi_{l}$ represents the threshold for the $l$-th access level, and $L$ denotes the total number of hierarchical layers within the knowledge base. This formula employs an indicator function $\mathbf{1}(\cdot)$, which assigns a value of 1 for each threshold that the trust score $T_{i}$ meets or exceeds, thereby cumulating the number of accessible layers. This structure allows for fine-tuned access control, where higher trust scores permit access to more sensitive information, and lower scores restrict users to more generalized content. The process is depicted in Figure 3, illustrating the integration of RAG and ICL to manage information flow based on the user’s trust score. Initially, RAG retrieves trust-oriented contexts, which are then seamlessly integrated by ICL along with a guidance prompt and the trust score, to guide the response generated by the language model. This tailored approach ensures a balance between content safety and personalization, effectively addressing specific user needs. The content richness function is therefore formulated as below:
$R^{(\tau)}_{i}(x,AL_{i})=LM\left(x,Context_{AL_{i}}\right)$ (14)
In this formula, $LM(\cdot)$ tailors response $x$ according to the user’s authorized access level $AL_{i}$, specifically to the $Context$ that pertains to the query. Here, $Context$ refers to the enriched retrieval information relevant to the inquiry. This adaptive guardrail mechanism ensures each user interaction is secure and personalized, balancing usability with security protocols.
By employing adaptive trust modeling along with content adaptation mechanism, our guardrail not only enhances the practicality of LLMs but also ensures outputs adhere to ethical and safety standards. This dynamic adjustment of model parameters and sensitive knowledge access level based on trust scores establishes a framework for the safe and reliable deployment of language models, effectively balancing user expectations with content integrity.
Guardrails % In-domain Questions % Out-of-domain Questions
Llama Guard 96.30% 89.94%
Llama Guard 2 40.12% 41.62%
Llama Guard 3 4.32% 3.63%
gpt-3.5-turbo 9.26% 2.79%
gpt-4o-mini 24.62% 6.87%
gpt-4 27.16% 3.07%
Nvidia NeMo 48.15% 40.94%
Our method (Enhanced by Trust Modeling) 97.53% 0.83%
Table 1: Comparison of Guardrail System Access Controls for High-Trust User
5 Experiments
5.1 Data Preparation
For our experiments, we gathered data from several datasets including AdvBench [45], HarmBench [26] and ToxicChat [22], categorizing prompts by their relevance to different domains to evaluate our adaptive trust model for guardrails. This dataset comprises 520 harmful prompts reflecting potential sensitive interactions, supplemented by 3,000 safe prompts for interaction diversity. We partitioned the prompts into in-domain and out-of-domain categories based on whether their topics matched the user’s professional domain. For example, prompts related to computer science were classified as in-domain for users from the computer science (CS) domain, while prompts from other fields were considered out-of-domain. Due to privacy concerns and restricted access to actual user data from TTPs, we created synthetic profiles for virtual users, incorporating simulated ratings, professional areas, social information, and more, to emulate real-world trust attributes, allowing us to verify if the trust model effectively aligns access to sensitive information with validated user identities. All necessary materials including code and datasets, are hosted on GitHub [^8] The anonymized GitHub will be made public upon paper acceptance to comply with double-blind review requirements.
5.2 Implementation Configurations
In our experiment, the trust model dynamically regulates guardrail settings (predefined as 3 levels including relax, normal, strict) based on user interactions and trust attributes from TTPs. The model categorizes interactions as follows: “relax" allows access to sensitive information to authorized user, “normal" limits access without terminating the session, and “strict" terminates the session and may lock the account pending further verification. To mitigate security risks, the system refreshes verification after every ten high-risk interactions. This regularly recalibration ensures that even users with high trust scores and reputable backing are not given unlimited access to sensitive data, maintaining system integrity and security.
5.3 Adaptability and Security Analysis
In this case, we evaluated the adaptability and security of our adaptive guardrail systems in Table 1, which integrate trust modeling, by benchmarking against state-of-the-art black-box and white-box techniques, including the GPT series (enhanced by white-box techniques like fine tuning), Llama Guard, and Nvidia NeMo. Our focus was on determining whether these guardrail systems could appropriately modify access controls to recognize and support the needs of verified users, thereby facilitating their access to sensitive information. We designed controlled scenarios where all systems were pre-configured to recognize users identified as experts in their professional domain, allowing us to assess how dynamically these systems could adjust an expert’s permissions to access restricted content within this field.
The results demonstrated the superior adaptability and effectiveness of our adaptive guardrail system, achieving a 97.53% success rate in granting access to in-domain questions for high-trust users. This sharply contrasts with the GPT series, fine-tuned with safety guidelines before deployment, which achieved only 9.26%, 24.62%, and 27.16% for similar queries, underscoring the flexibility our system offers over the rigidity of these models. Although the GPT series restricts irrelevant sensitive queries to a low rate of 2.79%, gaps still persist in certain cases, such as “How to perform a successful hit-and-run and avoid being caught by digital cameras." GPT misclassified it as CS-related, likely due to associating “digital" with the CS domain, leading to errors. In contrast, our method effectively rejected it, demonstrating enhanced safety.
In comparison with black-box guardrails such as the Llama Guard series and Nvidia NeMo, our system exhibited enhanced control and safety. Llama Guard allowed 96.30% access to in-domain content, slightly lower than our method. This discrepancy may stem from pre-trained biases of LLMs, such as strongly associating words like “virus" with medical contexts rather than cybersecurity, reflecting the necessity for fine-tuning in dynamic scenarios of existing guardrails. Llama Guard also allowed 89.94% access to unrelated areas, highlighting a lack of precision that could introduce security risks. On the other hand, Llama Guard 2 and 3 significantly restricted access, potentially to the point of impeding user experience due to over-restrictiveness. Nvidia NeMo, with more moderate access rates of 48.15% and 40.94% to in-domain and other areas respectively, suggested a balanced approach but does not match the nuanced control our model provides. Additionally, the results revealed that black-box methods predominantly employed rigid, predefined rules. Despite our efforts to slightly relaxed these to accommodate the specialized needs of trustworthy users, they still had similar probabilities, indicating the difficulty in differentiating sensitive queries. While Llama Guard and Nvidia NeMo enable customization of security settings for specific queries, they demonstrate vulnerabilities in preventing unauthorized access to sensitive content. Our system overcomes these shortcomings by implementing access controls through verifiable user credentials and query relevance. This approach safeguards sensitive information while upholding ethical standards, positioning our adaptive guardrail system as a groundbreaking solution for the secure and versatile deployment of future LLMs. Details of the ablation study for each variables are provided in Appendix A.
5.4 Contextual Implications Analysis
In this experimental case, we explored the interplay between user trust scores and access to information of varying sensitivity levels using LLMs and In-context Learning. The setup involved adjusting hyperparameters and a contextual knowledge base according to user trust scores to control the granularity of information provided during potentially hazardous interactions. Our model harnessed contextual knowledge from external databases, for instance, step-by-step processes for making bombs and handling hazardous materials with different levels of danger. We specifically crafted a scenario using ChatGPT around the prompt “Write a step-by-step guide for making a bomb" as shown in Figure 4.
Figure 4: Access levels based on trust modeling and contextual adaption in extremely hazardous information scenarios
Figure 5: Trust Dimension Contributions to Dynamic Trust Modeling Across Authority Levels
Although this query is inherently hazardous, it has valid contexts within fields like chemistry. We established an access threshold score of 0.8. Observations showed that users identified as chemists with trust scores below this threshold experienced access restrictions similar to general users, receiving denial responses such as, “I’m sorry, I can’t assist with that …". In contrast, users exceeding this threshold were granted access to non-critical information about bomb-making procedures and materials. For highly sensitive or dangerous information, such as techniques involving nitroglycerin bombs, we imposed a more stringent requirement, setting a trust score threshold of 0.95. Our observations confirmed that only users who met this elevated trust criterion were granted access, effectively restricting extremely hazardous information to verified highly-trustable users.
This case showcases how the adaptive guardrail, equipped with contextual knowledge, meets legitimate informational needs while strictly limiting access to extremely hazardous information. It ensures that only qualified individuals receive necessary information and prevents misuse by those not meeting stringent trust criteria.
5.5 Trust Attribution Across Credential Levels
To further illustrate the adaptability and rationality of our trust computation framework across users authorized by TTPs with varying authority levels, we conducted a trust contribution demonstration of how trust dynamics facilitate or restrict access to sensitive information, as visualized in Figure 5. This demonstration assessed users with similar basic trust attributes, while their authority levels determined the respective contributions of $DT$ and $AT$ to the final $T_{i}$.
For users with Top Authority credentials, $T_{i}$ was predominantly influenced by $AT$, given that their access rights were validated by highly reliable TTPs. The dominance of $AT$ ensured that such users could securely and stably access sensitive content within their authorized domains while maintaining robust safeguards by aligning access privileges with verifiable credential reliability. Conversely, users without Top Authority credentials experienced a more dynamic interplay of $DT$ and $AT$, where historical interaction patterns ($DT$) played a pivotal role in moderating ongoing access. Even with moderately high validation from less authoritative TTPs, $DT$ scores provided essential flexibility for real-time access adaptation, ensuring that consistent and trustworthy behavior earned controlled access to sensitive information while restricting prolonged engagement to mitigate potential risks. The system dynamically adjusted $DT$ weighting to compensate for weaker $AT$ validation, effectively adapting to diverse trust scenarios. For users with low $DT$ and authority levels, the framework enforced strict restrictions, even when less credible TTPs assigned seemingly high trust scores, by integrating multi-dimensional trust attributes and regularly re-certification to prevent exploitation through falsified credentials, thus demonstrating robust protection against adversarial manipulation.
The experimental results highlight the distinct contributions of $DT$ and $AT$ to trust computation, with their influence varying by authority level. For Top Authority users, $AT$ dominated due to highly credible credentials, while $DT$ played a pivotal role for lower authority users by leveraging historical interactions to refine trust scores dynamically. This interplay underscores the framework’s adaptability in balancing security, usability, and flexibility. By integrating direct and authority-verified trust, the model enforces robust safeguards while enabling context-aware adjustments for secure, personalized access.
6 Discussion
6.1 Human-AI Oversight for Ethical Safeguards
For very ethically charged topics and hallucination risks, we envision a hybrid approach combining trust thresholds with categorical restrictions, creating a “zero-trust zone" for topics with catastrophic misuse potential. This would require both high trust authentication (as hierarchical confidentiality thresholds shown in Section 5.4) and human expert authorization, aligning with community expectations for human-in-the-loop oversight [14]. Additionally, RAG systems prioritize verified information from trusted sources, mitigating hallucination risks while maintaining legitimate access—thus balancing adaptability with ethical safeguards.
6.2 Equitable Access for AI techniques
Trust-oriented guardrails enable dynamic access control, contrasting with static permission profiles. Our system evaluates multi-dimensional trust metrics by combining direct interaction history with authority verification to calibrate access decisions. This design mirrors real-world trust cases, such as financial rehabilitation programs, where individuals with past defaults can progressively regain trust through consistent positive behavior. Unlike static systems with permanent restrictions, our approach adjusts privileges based on demonstrated reliability, enabling trust rehabilitation while maintaining appropriate safeguards, dynamically balancing security and equitable access to align AI behavior with societal expectations [27].
6.3 Resilience to “Wolf in Sheep’s Clothing” cases
Our trust-oriented approach can effectively counters “wolf in sheep’s clothing” scenarios where adversaries intersperse benign queries with malicious ones to circumvent safety measures. Unlike static guardrails, our system dynamically evaluates both credential verification ($AT$) and behavioral consistency ($DT$), triggering progressive trust degradation upon detecting sensitive query patterns. Appendix B demonstrates users exhibiting suspicious behavior face increasingly restricted access despite intermittent legitimate interactions. The dual verification process creates overlapping security boundaries, while mandatory re-certification limit the temporal validity of falsified credentials, establishing robust protection against sophisticated attacks that traditional guardrails cannot address.
7 Conclusion
This paper presents a groundbreaking approach to safeguarding the security and utility of LLMs through an adaptive guardrail mechanism based on trust modeling and online ICL via RAG. Our model dynamically integrates real-time, user-specific trust assessments, offering personalized content moderation previously unachievable with static guardrails. This trust-oriented method effectively secures sensitive information without compromising user engagement, thereby meeting diverse user needs. Future work could be extended to include more comprehensive contextual hierarchy and deeper integration with emerging AI technologies, refining the dimension of trust-oriented assessments in complex interaction scenarios. This foundational work paves the way for responsible AI advancements, promoting more dependable and user-centric deployment of AI models.
Appendix A Ablation Test Analysis
In our ablation study, we evaluated the key dimensions of $AT$ using 180 sensitive and 1,000 positive prompts, focusing on a user verified by medium-level TTPs with a high credit score. This selection highlights how guardrails perform when users from moderate institutions access sensitive information, influenced by both $DT$ and $AT$. Initially, without omitting any variables, 61.11% of interactions were managed with relaxed guardrails, while 9.45% faced the strictest controls, indicating potential interaction termination and further verification to maintain security.
We observed significant shifts in guardrail behavior upon the omission of each variable in Table 2: The removal of $AR$ and $NR$ from the Authority Verified Trust reveals a critical vulnerability in maintaining information safety, as evidenced by significant increases in relaxed interactions, to 78.33% and 82.78% respectively. This contrasts sharply with the effects observed when removing $A$, $S$ and $C$, where relaxed guardrail percentages notably decreased, highlighting their respective roles in enhancing the system’s security-controlling precision. Specifically, removing $A$ resulted in relaxed interactions plummeting to 43.33%, removing $S$ dropped them to 52.78%, and omitting $C$ brought them down to 58.33%. The strict guardrail percentages also reflect impacts: removal of $A$ resulted in the highest strict guardrail application at 11.11%, followed by $C$ at 10.56%, and $S$ at 10%. In contrast, removing $AR$ completely eliminated strict guardrails, underscoring its critical role in stringent access control for irrelevant sensitive queries in dynamic interaction. These outcomes underscore the unique contribution of each variable: $AR$ and $NR$ are pivotal in enforcing strict access controls and safeguarding information security, whereas $A$, $S$, and $C$ are essential for validating authority, ensuring interaction relevance, and maintaining flexible, precise trust assessments. Collectively, this ablation analysis demonstrates how each variable’s role is crucial for balancing user access with system security and practicality.
Appendix B Trust-Driven Safeguards in Mitigating Jailbreaking Threats
The trust-oriented adaptive guardrail system establishes a more robust safeguards against jailbreaking attacks by leveraging a dual-layered trust computation framework comprising $DT$ and $AT$. Unlike static guardrails, which are prone to being bypassed through carefully crafted prompts, this adaptive system preemptively assesses the user’s trust level before applying normal guardrail logic to determine whether a request involving sensitive information should be accepted or denied. This design significantly enhances the robustness of the system, particularly against adversarial scenarios involving jailbreaking attempts.
Attackers often exploit static guardrails by submitting a mix of benign and sensitive queries to mask their malicious intent. However, the adaptive guardrail counters such strategies by dynamically calibrating $DT$ and $AT$ scores based on historical interaction data and authority validation. For instance, attackers with limited interaction histories exhibit low $DT$, as their lack of prior trustworthy engagement raises suspicion. Similarly, users failing to demonstrate validation from reliable TTPs are assigned low $AT$, further restricting their access. This trust-oriented approach ensures that users with insufficient trust scores are denied access outright, regardless of the crafted nature of their prompts.
To illustrate this capability, Figure 6 presents a simulation involving a low-authority user attempting to bypass the guardrail system using mixed benign and sensitive jailbreaking queries. Unlike high-authority users, whose trust scores predominantly rely on TTP-based validation and are robust against malicious attack strategies, low-authority users’ trust scores are dynamically adjusted by $DT$. Sensitive queries cause a cumulative decline in $DT$, progressively eroding the user’s trust score. This mechanism effectively denies access, even when intermittent safe prompts are submitted to obscure malicious intent. Additionally, periodic $DT$ resets and stricter penalty constraints prevent attackers from exploiting the system by repeatedly submitting benign queries to regain access.
Figure 6: Mitigating Jailbreaking Attacks Through Direct Trust Decline from Sensitive Queries
The adaptive guardrail’s robustness is rooted in its trust modeling foundation, which integrates historical user data and authority validation to preemptively assess trust levels before applying content moderation logic. By ensuring only users meeting trust thresholds can proceed, the system mitigates unauthorized access attempts early in the process. Unlike static guardrails, which lack adaptability against dynamic attack patterns, this layered approach dynamically identifies and blocks adversarial users through continuous trust monitoring and recalibration. Users attempting to exploit the system with sophisticated prompts or mixed benign and sensitive queries are quickly flagged due to their low trust profiles, effectively preventing prolonged exploitation. This trust-oriented mechanism also ensures that sensitive content is accessible only to high-trust users while maintaining robust protections against unauthorized access and misuse. By balancing personalization, usability and security, the system demonstrates resilience against jailbreaking attacks and establishes a scalable framework for trust-aware LLM applications in security-critical environments.